# Low-Resourced Speech Recognition for Iu Mien Language via Weakly-Supervised Phoneme-based Multilingual Pre-training

### v1

@@ 1 Introduction
Iu Mien word | syllable | initial consonant | vowel | tone | ||
initial | main | final | ||||
aanwatv | aan | - | - | aa | n | - |
watv | w | - | a | t | v | |
baengh | baengh | b | - | ae | ng | h |
nqaang | nqaang | nq | - | aa | ng | - |
guinh | guinh | g | ui | - | n | h |
@@ 4 Method
We add a linear layer on top of the acoustic encoder, which is used to map the feature vectors of the encoder output to the target prediction space. For methods of fine-tuning based on subwords, we use the random initialization method to initialize the linear layer parameters. For the phoneme-based approach to fine-tuning, in order to share the parameters from the pre-training as much as possible, we used the method in Whistle to use the parameters of the linear layer from the pre-training backbone for initialization. For the k-th row vector in the linear layer weight matrix at the top of the encoder during pre-training, it is considered as the embedding vector corresponding to the k-th phoneme in the predicted phoneme list. For a target language phoneme that has appeared in the predicted phoneme lexicon at pre-training, we use the embedding vector of that phoneme obtained at pre-training to initialize the corresponding position of the linear layer weight matrix at fine-tuning.
@@ 5.1 Dataset
MightLJSpeech is a publicly available dataset of the Iu Mien language. The dataset is 9.7 hours long, contains 9761 audios and corresponding Iu Mien language text annotations, and the audio file sampling rate is 22050Hz.
In the experiment, all audio files were first resampled to 16KHz, and then the dataset was divided into training set, validation set and test set in a ratio of 8:1:1. To test the statistical significance of our experimental results, we conducted each experiment three independent times and used the average of these independent runs as the final result. For each run, we divided the dataset using a cross-validation approach. Specifically, we split the original dataset into 10 parts. In each run, two parts were used as the development set and test set, while the remaining eight parts were used as the training set. The development set and test set selected for each run were not reused in other runs for development or test.
@@ 6.1 Results for subword-based FT
Table 2 shows the results for subword-based FT, starting from different pre-trained models. O1 denotes the subword-based monolingual baseline, i.e., trained from scratch. M1 denotes the result from fine-tuning of Whistle-small. M3 denotes the result from fine-tuning of a Wav2Vec2-base model, trained from scratch using self-supervised pre-training with the same training data as the Whistle-small model, which we call the Wav2Vec2-cv10 model. Wav2Vec2-cv10 was trained using the fairseq toolkit, following the wav2vec 2.0 base pre-training configuration provided with the toolkit. M4 denotes the result from fine-tuning of a subword pretraining model, called Mul10-subword, which uses the same pretraining data and encoder architecture as Whistle-small, but is modeled using subwords with a vocabulary size of 5000. To show the statistical significance of the experimental results, we conducted three independent cross-validation tests for each experiment and took the average as the final experimental result.

### v2

@@ 1 Introduction
Iu Mien word | Chinese meaning | Chinese meaning | syllable | initial consonant | vowel | tone | ||
in English | initial | main | final | |||||
ginghgungv | 蜻蜓 | dragonfly | gingh | g | - | i | ng | h |
gungv | g | - | u | ng | v | |||
baengh | 平 | flat | baengh | b | - | ae | ng | h |
nqaang | 后来 | later | nqaang | nq | - | aa | ng | - |
guinh | 圈,转 | circle, turn | guinh | g | u | i | n | h |
@@ 4 Method
We add a linear layer on top of the acoustic encoder, which is used to map the feature vectors of the encoder output to the target prediction space. For methods of fine-tuning based on subwords, we use the random initialization method to initialize the linear layer parameters. For the phoneme-based approach to fine-tuning, in order to share the parameters from the pre-training as much as possible, we used the method in Whistle to use the parameters of the linear layer from the pre-training backbone for initialization. For the k-th row vector in the linear layer weight matrix at the top of the encoder during pre-training, it is considered as the embedding vector corresponding to the k-th phoneme in the predicted phoneme list [2]. For a target language phoneme that has appeared in the predicted phoneme lexicon at pre-training, we use the embedding vector of that phoneme obtained at pre-training to initialize the corresponding position of the linear layer weight matrix at fine-tuning.
@@ 5.1 Dataset
MightLJSpeech is a publicly available dataset of the Iu Mien language. The dataset is 9.7 hours long, consisting of 9761 utterance from a female speaker and the corresponding Iu Mien language text annotations. The sampling rate for audio files is 22.05kHz.
In the experiment, all audio files were first resampled to 16KHz, and then the dataset was divided into training set, development set and test set in a ratio of 8:1:1. To test the statistical significance of our experimental results, we conducted each experiment three independent times and used the average of these independent runs as the final result. For each run, we divided the dataset using a cross-validation approach. Specifically, we split the original dataset into 10 parts. In each run, two parts were used as the development set and test set, while the remaining eight parts were used as the training set. The development set and test set selected for each run were not reused in other runs for development or test.
@@ 6.1 Results for subword-based FT
Table 2 shows the results for subword-based FT, starting from different pre-trained models. O1 denotes the subword-based monolingual baseline, i.e., trained from scratch. M1 denotes the result from fine-tuning of Whistle-small. M3 denotes the result from fine-tuning of a Wav2Vec2-base model, trained from scratch using self-supervised pre-training with the same training data as the Whistle-small model, which we call the Wav2Vec2-cv10 model. Wav2Vec2-cv10 was trained using the fairseq toolkit, following the wav2vec 2.0 base pre-training configuration provided with the toolkit. M4 denotes the result from fine-tuning of a subword pretraining model, called Mul10-subword, which uses the same pretraining data and encoder architecture as Whistle-small, but is modeled using subwords with a vocabulary size of 5000. To fairly compare the performance of models trained with the same pretraining data using different methods, the parameter count of Wav2Vec2-cv10 and Mul10-subword is kept similar to Whistle-small (90M). To show the statistical significance of the experimental results, we conducted three independent cross-validation tests for each experiment and took the average as the final experimental result.